# Beyond Optimal Fault-Tolerance

### v3

@@ 1. Introduction
This paper provides a thorough investigation of this question. To expose the richness of the answer, we work with a timing model that can be viewed as an interpolation between the synchronous and partially synchronous settings. In addition to the usual parameters $\Delta$ and GST (known and unknown, respectively) of the partially synchronous model, we allow for a known parameter $\Delta^{*}\geq\Delta$ which may or may not bound message delays prior to GST. Canonically, $\Delta^{*}$ should be thought of as orders of magnitude larger than $\Delta$ (days or weeks, as opposed to millisecond or seconds), with $\Delta$ indicating the speed of communication between processes when all is well (no network issues, no attacks) and $\Delta^{*}$ the speed of (possibly out-of-band) communication around the time of an attack. [^1] Indeed, for our positive results, message delays must be bounded by $\Delta^{*}$ for the duration of our recovery procedure, but not otherwise. We will be interested in protocols that always satisfy all the “usual” guarantees (1) and that finalize transactions in time $O(\Delta)$ (rather than $O(\Delta^{*})$) after GST, whether or not pre-GST message delays are bounded by $\Delta^{*}$, and also satisfy the additional recovery guarantees (2) and (3) in the event that pre-GST message delays are in fact bounded by $\Delta^{*}$. [^2] In particular, a synchronous protocol with respect to the parameter $\Delta^{*}$ will not generally satisfy consistency and liveness if message delays do not happen to bounded above by $\Delta^{*}$.
@@
  • $\tfrac{1}{3}$-resilience in partial synchrony (independent of whether $\Delta^{*}$ bounds pre-GST message delays), with worst-case latency $O(f_{a}\Delta)$ after GST, where $f_{a}$ denotes the (actual) number of faulty processes;
  • should pre-GST message delays be bounded by $\Delta^{*}$, recovers from consistency violations in time $O(f_{a}\cdot\Delta^{*})$ and with rollback $2\Delta^{*}$; and
@@ 2. The setup
Message delays. We consider a discrete sequence of timeslots $t\in\mathbb{N}_{\geq 0}$. As discussed in the introduction, we consider protocols that operate in partial synchrony (with some parameter $\Delta$, perhaps in the order of seconds or milliseconds) and satisfy additional recovery properties should synchrony hold (with some different parameter $\Delta^{*}$, perhaps in the order of days).
@@
State machine replication. A protocol for state machine replication (SMR) must cause correct processes to finalize logs (sequences of transactions) that are live and consistent with each other. Formally, if $\sigma$ and $\tau$ are sequences, we write $\sigma\subseteq\tau$ to denote that $\sigma$ is a prefix of $\tau$. We say $\sigma$ and $\tau$ are compatible if $\sigma\subseteq\tau$ or $\tau\subseteq\sigma$. If two sequences are not compatible, they are incompatible. If $\sigma$ is a sequence of transactions, we write $\mathtt{tr}\in\sigma$ to denote that the transaction $\mathtt{tr}$ belongs to the sequence $\sigma$.
@@ 4. The intuition behind the wrapper
We describe a wrapper, which takes an accountable and optimally resilient SMR protocol $\mathcal{P}$ as input, and which runs an execution of $\mathcal{P}$ until a consistency violation occurs [^11] By “optimally resilient,” we mean that the protocol’s consistency resilience $\rho_{C}$ and liveness resilience $\lambda_{L}$ in partial synchrony are both positive and satisfy $\rho_{C}+2\rho_{L}=1$ (as is the case for all of the “usual” SMR protocols designed for the partially synchronous setting). This assumption is merely to simplify the presentation. For a non-optimally resilient protocol, the “$1-\rho_{L}$” term in (1) should be replaced by “$(1+\rho_{C})/2$”. Once this happens, the wrapper triggers a ‘recovery procedure’, which achieves consensus on a set of faulty processes $F$ for which a proof of guilt exists, together with a long initial segment of the log produced by $\mathcal{P}$ below which no consistency violation has occurred. The wrapper then initiates another execution of $\mathcal{P}$ that takes this log as its genesis log, with the players in $F$ removed from the process set. This next execution is run until another consistency violation occurs, and so on.
@@
The structure of this section. In Section 4.1, we describe the intuition behind a feature of the wrapper which allows us to ensure rollback bounded by $2\Delta^{*}$. Section 4.2 describes the intuition behind the recovery procedure.

### v4

@@ 1. Introduction
This paper provides a thorough investigation of this question. To expose the richness of the answer, we work with a timing model that can be viewed as an interpolation between the synchronous and partially synchronous settings. In addition to the usual parameters $\Delta$ and GST (known and unknown, respectively) of the partially synchronous model, we allow for a known parameter $\Delta^{*}\geq\Delta$ which may or may not bound message delays prior to GST. Canonically, $\Delta^{*}$ should be thought of as orders of magnitude larger than $\Delta$, with $\Delta$ indicating the speed of communication between processes when all is well (no network issues, no attacks) and $\Delta^{*}$ a safe (and possibly large) upper-bound on the speed of (possibly out-of-band) communication around the time of an attack. [^1] Indeed, for our positive results, message delays must be bounded by $\Delta^{*}$ for the duration of our recovery procedure, but not otherwise. We will be interested in protocols that always satisfy all the “usual” guarantees (1) and that finalize transactions in time $O(\Delta)$ (rather than $O(\Delta^{*})$) after GST, whether or not pre-GST message delays are bounded by $\Delta^{*}$, and also satisfy the additional recovery guarantees (2) and (3) in the event that pre-GST message delays are in fact bounded by $\Delta^{*}$. [^2] In particular, a synchronous protocol with respect to the parameter $\Delta^{*}$ will not generally satisfy consistency and liveness if message delays do not happen to bounded above by $\Delta^{*}$.
@@
  • $\tfrac{1}{3}$-resilience in partial synchrony (independent of whether $\Delta^{*}$ bounds pre-GST message delays), with expected latency $O(\Delta)$ after GST;
  • should pre-GST message delays be bounded by $\Delta^{*}$, recovers from consistency violations in expected time $O(\Delta^{*})$ and with rollback $2\Delta^{*}$; and
@@
A comment on the benefits of automated recovery. It is sometimes assumed in the literature (e.g. (Sridhar et al., 2023)) that, in the event of a consistency violation, an ‘administrator’ will remove perpetrators from the system and co-ordinate an appropriate ‘reboot’. Our automated recovery procedure has the significant benefit that it does not rely on a centralized entity and single point of failure (the very point of distributed systems) , and that it also formalizes a process by which one can guarantee bounded rollback (a principal focus of this paper).
@@ 2. The setup
Message delays. We consider a discrete sequence of timeslots $t\in\mathbb{N}_{\geq 0}$. As discussed in the introduction, we consider protocols that operate in partial synchrony (with some parameter $\Delta$, perhaps in the order of seconds or milliseconds) and satisfy additional recovery properties should synchrony hold (with some different parameter $\Delta^{*}$, which may be set much larger than $\Delta$) while running the ‘recovery procedure’.
@@
State machine replication. Informally, a protocol for state machine replication (SMR) must cause correct processes to finalize logs (sequences of transactions) that are live and consistent with each other, and must also produce ‘certificates’ of finalization that can be presented to clients who may not always be online and observing the protocol execution. Formally, if $\sigma$ and $\tau$ are sequences, we write $\sigma\subseteq\tau$ to denote that $\sigma$ is a prefix of $\tau$. We say $\sigma$ and $\tau$ are compatible if $\sigma\subseteq\tau$ or $\tau\subseteq\sigma$. If two sequences are not compatible, they are incompatible. If $\sigma$ is a sequence of transactions, we write $\mathtt{tr}\in\sigma$ to denote that the transaction $\mathtt{tr}$ belongs to the sequence $\sigma$.
@@ 4. The intuition behind the wrapper
We describe a wrapper, which takes an accountable and optimally resilient SMR protocol $\mathcal{P}$ as input, and which runs an execution of $\mathcal{P}$ until a consistency violation occurs. [^11] By “optimally resilient,” we mean that the protocol’s consistency resilience $\rho_{C}$ and liveness resilience $\lambda_{L}$ in partial synchrony are both positive and satisfy $\rho_{C}+2\rho_{L}=1$ (as is the case for all of the “usual” SMR protocols designed for the partially synchronous setting). This assumption is merely to simplify the presentation. For a non-optimally resilient protocol, the “$1-\rho_{L}$” term in (1) should be replaced by “$(1+\rho_{C})/2$”. Once this happens, the wrapper triggers a ‘recovery procedure’, which achieves consensus on a set of faulty processes $F$ for which a proof of guilt exists, together with a long initial segment of the log produced by $\mathcal{P}$ below which no consistency violation has occurred. The wrapper then initiates another execution of $\mathcal{P}$ that takes this log as its genesis log, with the players in $F$ removed from the process set. This next execution is run until another consistency violation occurs, and so on.
@@
The structure of this section. In Section 4.1, we describe the intuition behind a feature of the wrapper which allows us to ensure rollback bounded by $2\Delta^{*}$. We stress that bounding rollback is non-trivial: this is the requirement on the recovery procedure that requires the most delicate analysis. Section 4.2 then describes the intuition behind the recovery procedure.